# Understanding the Multi-modal Prompts of the Pre-trained Vision-Language Model

### 18 Dec 2023 v1

@@ Abstract
Prompt learning has emerged as an efficient alternative for fine-tuning foundational models, such as CLIP, for various downstream tasks. However, there is no work that provides a comprehensive explanation for the working mechanism of the multi-modal prompts. In this paper, we conduct a direct analysis of the multi-modal prompts by asking the following questions: $(i)$ How do the learned multi-modal prompts improve the recognition performance? $(ii)$ What do the multi-modal prompts learn? To answer these questions, we begin by isolating the component of the formula where the prompt influences the calculation of self-attention at each layer in two distinct ways, i.e., $(1)$ introducing prompt embeddings makes the $[cls]$ token focus on foreground objects. $(2)$ the prompts learn a bias term during the update of token embeddings, allowing the model to adapt to the target domain. Subsequently, we conduct extensive visualization and statistical experiments on the eleven diverse downstream recognition datasets. From the experiments, we reveal that the learned prompts improve the performance mainly through the second way, which acts as the dataset bias to improve the recognition performance of the pre-trained model on the corresponding dataset. Based on this finding, we propose the bias tuning way and demonstrate that directly incorporating the learnable bias outperforms the learnable prompts in the same parameter settings. In datasets with limited category information, i.e., EuroSAT, bias tuning surpasses prompt tuning by a large margin. With a deeper understanding of the multi-modal prompt, we hope our work can inspire new and solid research in this direction.
@@ 1 Introduction
In the alignment contribution statistic, we count the relevance between inputs and $[cls]$ token during alignment. In the statistics results of the text branch, we observe that the contribution of category tokens does not increase but decreases. On the contrary, the contribution of $t_{SOS}$ and prompt tokens significantly increased, especially the $t_{SOS}$ token. For the image branch, we reveal that the learned multi-modal prompts do not significantly adjust the relative relevance to $[cls]$ between inputs. Surprisingly, the contribution of multi-modal prompts to the alignment process is significant.
@@
Based on our findings, we propose a novel bias tuning that directly incorporates the learnable bias to each transformer block to validate the importance of the dataset bias. Comparative experiments demonstrate that bias tuning outperforms prompt tuning with the same number of parameters, particularly on datasets with limited category information, i.e., EuroSAT [14]. However, the drawbacks of directly learning bias are also evident, namely the limited design options for bias and the finite number of learnable parameters in each layer. In contrast, there is a wide variety of prompt options that can be added in each layer. To search for the optimal hyperparameter of the prompt for each dataset, we ablate on the number and layer of prompts. Bias tuning still falls slightly short compared to the optimal prompt configuration. However, for datasets with limited category information, the advantages of bias tuning remain highly evident.
Main Findings and Contributions: (1) We probe to understand the working mechanism of the multi-modal prompts for the pre-trained vision-language model, conducting a series of extensive experiments from various perspectives, including alignment relevance, attention, attention distance, attention entropy, and visualization. Our findings show that the prompts mainly learn as the dataset bias. (2) Based on the findings, we propose a novel bias tuning that directly incorporates the learnable bias to each transformer block. Our bias tuning outperforms prompt tuning with the same number of parameters, particularly on datasets with limited category information, i.e., EuroSAT [14].
@@ 2.1 Vision-Language Pre-trained Model (VLM)
Figure 2: Statistics results on the contribution of each token of the language branch to the alignment task for the OxfordPet dataset. The corresponding two realizations are averaged over the statistics for the entire dataset, where the blue and red regions represent the statistics for zero-shot and prompt tuning, respectively. On the left side of the red vertical line, we show the contribution of the $\bm{t}_{SOS}$ token followed by four learned text prompts $\bm{P}_{t}$. On the right side, we show the $\bm{c}_{k}$ tokens and padding tokens. The horizontal coordinate is the index of the token and the vertical coordinate is the relative value of the contribution. For the zero-shot clip, we use the temple: ‘‘a photo of [Category].’’. For the tokens after the $\bm{t}_{EOS}$, we set their contribution as 0 for intuitive comparison. Meanwhile, we zoom in on comparing the index at which the prompt is placed in the middle of each figure. The statistics results of other datasets are detailedly shown in the Appendix.
Figure 3: Statistics results on the contribution of each patch of the image branch to the alignment task for the OxfordPet dataset. The corresponding two realizations are averaged over the statistics for the entire dataset, where the blue and red regions represent the statistics for zero-shot and prompt tuning, respectively. On the left side of the red vertical line, we show the contribution of the input patch tokens $\bm{\tilde{X}_{p}}$. On the right side, we show the learned vision prompts $\bm{P}_{v}$. The horizontal coordinate is the index of the token and the vertical coordinate is the relative value of the contribution. For the zero-shot clip, the values of the 196th-199th index are empty due to the absence of the vision prompts. Meanwhile, we zoom in on comparing the index at which the prompt is placed in the middle of each figure. The statistics results of other datasets are detailedly shown in the Appendix.
@@ 3.2.2 Aligning Contribution Statistics
To visually illustrate the dependency of alignment tasks on image patches and text tokens after incorporating prompts, we conduct a statistical analysis on eleven fine-grained recognition datasets. Before visualization, we preprocess the contribution according to normalization for intuitive comparison. The distribution of relevance for the OxfordPet dataset is presented in Fig.2 and Fig.3 (Ten more other datasets are shown in the Appendix).
  • In the statistics results of the text branch, we observe that the contribution of category tokens does not increase but decreases. On the contrary, the contribution of $t_{SOS}$ and prompt tokens significantly increased, especially the $t_{SOS}$ token.
@@ 3.4 Validation for the importance of the bias
To validate the conclusion, we directly add the learnable bias to the pre-trained model and compare it with the prompt tuning. In the bias tuning setting, we add an independent learnable embedding and repeat it to add to all tokens after each transformer layer of the image and language branch. After learning in the same few-shot setting, we compare the bias tuning, prompt tuning, and zero-shot CLIP in Table.1. The results show that bias tuning outperforms prompt tuning, especially on datasets with single-category information and scenes, such as EuroSAT [14]. However, prompt tuning has an advantage over bias tuning in that their performance can improve with an increase in the number of prompts. Table.1 shows the slight performance gap (0.5) between bias tuning and the optimal prompt tuning settings (acquired from the sensitivity analysis results in the Appendix) still exists.
Table 1: Validating the importance of bias. We compare bias with prompt tuning in the same parameter and the optimal settings.
Dataset CLIP PROMPT BIAS OPTIMAL
ImageNet [8] 66.7 72.5 71.9 73.1
Caltech101 [11] 93.3 95.2 96.0 95.2
DTD [7] 44.1 68.8 69.6 70.5
EuroSAT [14] 48.4 71.7 86.5 82.6
FGVCAircraft [24] 24.9 37.7 41.0 40.9
Food101 [2] 85.9 87.3 86.0 87.1
OxfordFlowers [26] 70.7 94.7 94.7 96.5
OxfordPets [27] 89.1 93.6 92.7 93.1
StanfordCars [20] 65.6 76.6 77.7 80.5
SUN397 [36] 62.6 75.3 74.7 76.2
UCF101 [33] 67.5 82.1 83.5 84.2
AVERAGE 65.3 77.8 79.5 80.0
3.5 Sensitivity Analysis for Multi-Modal Prompts
We analyze the robustness and sensitivity of the multi-modal prompts. The detailed results are shown in the Appendix.
Prompt Layer: We observe that the performance of the multi-modal prompts increases as the number of layers increases, but the upward trend also gradually slows down. Meanwhile, we observe that different datasets have different optimal numbers of layers. For datasets with many categories, such as ImageNet [8], the performance continues to increase with more layers. However, for datasets with fewer categories and high similarity between categories, such as EuroSAT [14], the performance saturates after adding a smaller number of multi-modal prompt layers.
Prompt Number:Increasing the number of prompts can improve performance in the pre-trained model. However, when the number becomes excessive, the stability of the learning process decreases. We observe that when the number of multi-modal prompts exceeds 20, the learning process of the prompts becomes highly unstable and prone to failure, especially for datasets with limited category information such as EuroSAT [14].
@@ 4 Conclusion
This paper probes to understand the working mechanism of the multi-modal prompts for the pre-trained vision-language model, conducting a series of extensive experiments from various perspectives, including alignment relevance, attention, attention distance, attention entropy, and visualization. Our findings show that the prompts mainly learn as the dataset bias. Based on the findings, we propose a novel bias tuning that directly incorporates the learnable bias to each transformer block that outperforms prompt tuning with the same number of parameters, particularly on datasets with limited category information.
@@ Acknowledgment
Appendix A Related Works
A.1 Vision Language models
A.2 Prompt learning
Prompt learning is an alternative fine-tuning method for transferring a model towards downstream tasks without re-learning the trained model parameters. This approach adapts a pre-trained model by adding a small number of new learnable embeddings at the input known as prompt tokens. Due to its efficiency in terms of parameters and convergence rate, prompt learning is found to be of great interest for adapting foundational models like CLIP for vision [17, 42, 34, 35] and vision-language tasks [44, 43, 46, 9]. CoOp [44] fine-tunes CLIP by optimizing a continuous set of prompt vectors in its language branch for few-shot image recognition. Bahng et al. [1] perform visual prompt tuning on CLIP by learning prompts on the vision branch. [6] and [22] propose to learn multiple sets of prompts for learning different contextual representations. CoCoOp [43] highlights the overfitting problem of CoOp and proposes to condition prompts based on visual features for improved performance on generalization tasks. MaPLe [18] proposes a multi-modal prompt learning approach by learning hierarchical prompts jointly at the vision and language branches of CLIP for better transfer. Our approach builds on a variant [31] where prompts are learned at both the vision and language encoder of CLIP.
Appendix B Detailed Formulation
Adding the multi-modal prompts, which learn from the few-shot training, often results in substantial recognition performance gains. Although the multi-modal prompts serve as tokens similar to the input tokens, it is also not rational to intuitively conclude how they work, since they are all at the high latitude abstraction space level, and we cannot determine what kind of image or text is learned. For the vision/text encoder, which consists of self-attention and MLP block, we analyze the impact of the added prompts in turn. For the MLP that attributes with independent operations for inputs, the added prompts through concatenate operations do not affect the upgrade of the input tokens. For self-attention layers, the prompts affect the attention weights of the $[cls]$ and input tokens in the attention calculation. We take the visual branch as an example to derive, after adding the prompt token ${X}$rightarrow ${X_{p}}=[X,P]$ (in the image branch ${x_{0}}$ denotes ${x_{cls}}$). The detailed self-attention calculation process is shown as follows:
Before adding the prompts:
$\Delta x_{i}$ $=\operatorname{softmax}\left(\frac{x_{i}\cdot Q\cdot(X\cdot K)^{T}}{\sqrt{d}}\right)\cdot X\cdot V$ (17)
[…]
$=\operatorname{softmax}\left(\frac{x_{i}\cdot Q\cdot K^{T}\cdot X^{T}}{\sqrt{d}}\right)\cdot X\cdot V$ (19)
$\quad\quad\quad\quad\quad\downarrow A\triangleq\frac{Q\cdot K^{T}}{\sqrt{d}}$ (20)
$=\operatorname{softmax}\left(x_{i}\cdot A\cdot X^{T}\right)\cdot X\cdot V$ (21)
[…]
$=\frac{\sum_{c=0}^{n}\left(e^{x_{i}Ax_{c}^{T}}\cdot x_{c}V\right)}{\sum_{c=0}^{n}e^{x_{i}Ax_{c}^{T}}}$ (23)
After Adding the prompts:
$\Delta x_{i}$ $=\operatorname{softmax}\left(\frac{x_{i}\cdot Q\cdot([X,P]\cdot K)^{T}}{\sqrt{d}}\right)\cdot[X,P]\cdot V$ (24)
[…]
$=\operatorname{softmax}\left(\frac{x_{i}\cdot Q\cdot K^{T}\cdot\left[X^{T},P^{T}\right]}{\sqrt{d}}\right)\cdot[X,P]\cdot V$ (26)
$\quad\quad\quad\quad\quad\quad\quad\quad\downarrow A\triangleq\frac{Q\cdot K^{T}}{\sqrt{d}}$ (27)
$=\operatorname{softmax}\left(x_{i}\cdot A\cdot\left[X^{T},P^{T}\right]\right)\cdot[X,P]\cdot V$ (28)
[…]
$=\frac{\sum_{c=0}^{n}\left(e^{x_{i}Ax_{c}^{T}}\cdot x_{c}V\right)+\sum_{c=0}^{num_{p}}\left(e^{x_{i}Ap_{c}^{T}}\cdot p_{c}V\right)}{\sum_{c=0}^{n}e^{x_{i}Ax_{c}^{T}}+\sum_{c=0}^{nump}e^{x_{i}Ap_{c}^{T}}}$ (30)
[…]
$=\frac{\sum_{c=0}^{n}\left(e^{x_{i}Ax_{c}^{T}}\cdot x_{c}V\right)}{\sum_{c=0}^{n}e^{x_{i}Ax_{c}^{T}}+\sum_{c=0}^{num_{p}}e^{x_{i}Ap_{c}^{T}}}+\frac{\sum_{c=0}^{num_{p}}\left(e^{x_{i}Ap_{c}^{T}}\cdot p_{c}V\right)}{\sum_{c=0}^{n}e^{x_{i}Ax_{c}^{T}}+\sum_{c=0}^{num_{p}}e^{x_{i}Ap_{c}^{T}}}$ (32)
In order to intuitively display the changes in attention before and after adding the prompt, we can simplify the formula as shown below where we simplify and omit the frozen parameters.
$x_{i}^{\ell+1}$ $=x_{i}^{\ell}+\sum_{j=0}^{M}\underline{\underline{a_{ij}}}\cdot x_{j},$ (33)
[…]
$x_{i}^{\ell+1}$ $=x_{i}^{\ell}+\sum_{j=0}^{M}\underline{\underline{a_{ij}^{\prime}}}\cdot x_{j}+\underline{\underline{\sum_{k=0}^{V}a_{ik}^{\prime}\cdot p_{v}^{k}}},$
Where ${a_{ij}},{a^{\prime}_{ij}}\propto{e^{x_{i}Ax_{j}}}$, ${a^{\prime}_{ik}}\propto{e^{x_{i}Ap^{k}_{v}}}$ and $A=\frac{{QK}^{T}}{{\sqrt{d}}}$, ${Q}$, ${K}\in\mathbb{R}^{{D_{v}}\times{D_{v}}}$ are frozen projection matrix. For the deep independent multi-modal prompts, the prompts are independently learned at each layer; once learned, they become a fixed piece of information for all data in the dataset, similar to the concept of bias. However, we observe that the information provided by the prompts is closely related to the patch tokens. That is, the more similar $x_{i}$ is to $p_{c}$, the larger the attention value obtained. Therefore, the information obtained by tokens from prompts in each layer differs. This results in changes to the values of the tokens in the subsequent layers, and the magnitude of these changes varies. As a result, the relative attention in the subsequent layers between tokens also changes. Thus, we make the following hypothesis about the working mechanism.
  • $\bullet$
  • $\bullet$
  • $\bullet$
    Both of the above are indispensable.
Table 2: Ablating the number of the multi-modal prompt layers. In each layer, the number of added multi-modal prompts is set to 8 which contains 4 vision prompts and 4 text prompts. The results are the average over seeds 1, 2, 3. More detailed results for each dataset and each seed are shown in the Appendix.
Method CLIP[28] Multi-Modal Prompt Layer
1 2 3 4 5 6 7 8 9 10 11 12
ImageNet [8] 66.70 71.70 71.73 71.80 71.93 72.17 72.20 72.43 72.57 72.50 72.77 72.83 72.90
Caltech101 [11] 93.30 94.93 95.17 95.17 95.43 95.20 95.10 95.53 95.17 95.37 95.37 94.97 95.93
DTD [7] 44.10 62.13 64.20 65.07 64.97 65.00 66.13 67.47 67.93 67.77 68.53 68.53 69.20
EuroSAT [14] 48.40 48.53 67.07 73.53 78.33 77.17 78.37 79.67 82.50 82.30 82.87 81.83 77.87
FGVCAircraft [24] 24.90 34.23 32.93 34.17 35.83 35.97 36.27 37.20 38.40 38.87 38.93 39.93 39.37
Food101 [2] 85.90 87.17 86.93 87.37 87.07 87.30 87.23 87.33 87.30 87.33 87.23 87.40 87.23
OxfordFlowers [26] 70.70 91.83 92.77 92.93 94.50 94.47 95.07 95.27 95.60 96.00 96.00 96.03 95.83
OxfordPets [27] 89.10 92.57 93.03 93.13 93.00 93.50 93.37 93.53 93.43 93.43 93.73 93.70 93.33
StanfordCars [20] 65.60 73.70 74.63 75.57 75.87 76.27 77.23 78.20 78.10 79.07 79.13 79.30 78.60
SUN397 [36] 62.60 73.80 74.00 74.30 74.43 74.77 75.47 75.17 75.93 76.00 76.00 75.87 75.93
UCF101 [33] 67.50 78.40 79.33 80.17 80.27 80.97 81.83 81.67 82.17 83.03 82.53 82.70 83.33
Average 65.35 73.54 75.62 76.66 77.42 77.53 78.02 78.50 79.01 79.24 79.37 79.37 79.05
Appendix C Sensitivity Analysis for Multi-Modal Prompts
In this section, we analyze the robustness and sensitivity of the multi-modal prompts by examining the number of multi-modal prompt layers and multi-modal prompts. For the ablation experiments, we show the average results over three seeds, i.e., seeds 1, 2, 3.
C.1 Ablating the number of the prompt layers
For ablating the multi-modal prompt layers, we conduct an experiment over 11 recognition datasets  [8, 11, 7, 14, 24, 2, 26, 27, 20, 36, 33] where the performance on each dataset is compared by adding different layers of multi-modal prompts and showing the results in Table.2. In each layer, the number of added multi-modal prompts is set to 8, which contains 4 vision prompts and 4 text prompts. From the results in the table, we observe that the performance of the multi-modal prompts increases as the number of layers increases, but the upward trend also gradually slows down. Meanwhile, we observe that different datasets have different optimal numbers of layers. For datasets with a large number of categories, such as ImageNet [8], the performance continues to increase with the addition of more layers. However, for datasets with fewer categories and high similarity between categories, such as EuroSAT [14], the performance saturates after adding a smaller number of multi-modal prompt layers. With the addition of the 10 and 11-layer multi-modal prompts, the average performance reached its highest, exceeding CLIP by more than 14 points.
C.2 Ablating the number of prompts
In table.3, we ablate the number of multi-modal prompts. We set the number of layers to 12 and the number of text and vision prompts is the same for each layer. We set the number of multi-modal prompts to 2, 4, 6, 8, 20, 40, 100, i.e., the number of text and vision prompts is set to 1, 2, 3, 4, 10, 20, 50, respectively. We find that increasing the number of prompts can lead to improved performance in the pre-trained model. However, when the number becomes excessive, the stability of the learning process decreases. We observe that when the number of multi-modal prompts exceeds 20, the learning process of the prompts becomes highly unstable and prone to failure, especially for datasets with limited category information such as EuroSAT [14].
Table 3: Ablating the number of the multi-modal prompts. In the experiment, we set the number of layers to 12. The results are the average over seeds 1, 2, 3. More detailed results for each dataset and each seed are shown in the Appendix.
Method IN[8] C101[11] DTD[7] FGV[24] F101[2] Flowers[26] Pets[27] Cars[20] SUN397[36] UCF[33] Euro[14]
CLIP[28] 66.70 93.30 44.10 24.90 85.90 70.70 89.10 65.60 62.60 67.50 48.40
2 prompts 72.50 95.17 68.77 37.73 87.27 94.73 93.57 76.63 75.33 82.13 71.73
4 prompts 72.63 95.67 68.63 37.27 87.37 95.63 93.60 77.77 75.70 82.53 76.53
6 prompts 72.90 95.47 69.50 39.90 87.40 95.87 93.23 79.00 75.97 82.97 79.37
8 prompts 72.90 95.93 69.20 39.37 87.23 95.83 93.33 78.60 75.93 83.33 77.87
20 prompts 73.13 95.20 70.47 40.93 87.07 96.47 93.07 80.50 76.23 84.17 82.60
40 prompts 72.97 95.73 70.73 41.90 87.03 96.20 93.53 81.10 76.03 84.27 -
100 prompts 72.77 95.57 68.53 42.47 86.40 96.20 92.97 81.33 76.20 83.83 -
Appendix D Attention Entropy
Attention Entropy: We analyze the attention of the model before and after adding prompts, as shown in the Fig.7, for probing the impact of incorporating vision prompts on the vision encoder’s feature extraction mechanism. The entropy value of attention reflects the focus of attention. A higher entropy value indicates a more uniform attention distribution and a global focus of attention. A lower entropy value indicates a more extreme attention distribution and a local focus of attention. In addition to calculating the attention entropy of all tokens in each layer, we also computed the attention entropy of the $[cls]$ token with respect to other tokens. From the results, we obtained the same conclusion as the attention distance statistic results: the prompts do not significantly impact the attention scope.
Figure 7: The attention entropy in different attention heads (dots). For the CLIP, we statistic the attention entropy and $[cls]$ token attention entropy of the original CLIP model. For the W/O Prompt, we statistic the attention entropy and $[cls]$ token attention entropy of the CLIP model after adding the prompts, but we remove the attention of the prompts. For the W Prompt, the attention of the prompts is retained.
Appendix E Alignment Contribution Statistic for CLIP, Shallow and Deep Prompt
In this section, we provide a detailed comparison of the alignment contribution between CLIP, Shallow Prompts, and Deep Prompts on other datasets. To ensure a fair comparison, we only compare on datasets where the clip template is fixed as “a photo of a + {} + {category}”. Since the templates for the DTD and EuroSAT datasets are “{} texture” and “a center satellite photo of + {category}” respectively, we do not make the comparison. Because the zero-shot performance of CLIP for the “a photo of a + {} + {category}” temple prompt on these two datasets is very bad. In the shallow prompt settings, we only add learned prompts in the first layer of the text and image encoder. In the deep prompt settings, we add learned prompts in all layers. We observe that in most datasets, the contribution of category tokens decreases after adding the learned prompt, while the contribution of the learned prompt tokens increases compared to that of template tokens. This phenomenon can be clearly observed in both shallow and deep prompts.
Figure 8: Statistics results on the contribution of each patch of the language branch to the alignment task for the ImageNet [8] dataset.
Figure 8: Statistics results on the contribution of each patch of the language branch to the alignment task for the ImageNet [8] dataset.
Figure 9: Statistics results on the contribution of each patch of the image branch to the alignment task for the ImageNet [8] dataset.
Figure 10: Statistics results on the contribution of each patch of the language branch to the alignment task for the FgvcAircrafat [24] dataset.
Figure 10: Statistics results on the contribution of each patch of the language branch to the alignment task for the FgvcAircrafat [24] dataset.
Figure 11: Statistics results on the contribution of each patch of the image branch to the alignment task for the FgvcAircrafat [24] dataset.
Figure 12: Statistics results on the contribution of each patch of the language branch to the alignment task for the OxfordFlowers [26] dataset.
Figure 12: Statistics results on the contribution of each patch of the language branch to the alignment task for the OxfordFlowers [26] dataset.
Figure 13: Statistics results on the contribution of each patch of the image branch to the alignment task for the OxfordFlowers [26] dataset.
Figure 14: Statistics results on the contribution of each patch of the language branch to the alignment task for the StanfordCars [20] dataset.
Figure 14: Statistics results on the contribution of each patch of the language branch to the alignment task for the StanfordCars [20] dataset.
Figure 15: Statistics results on the contribution of each patch of the image branch to the alignment task for the StanfordCars [20] dataset.
Figure 16: Statistics results on the contribution of each patch of the language branch to the alignment task for the UCF101 [33] dataset.
Figure 16: Statistics results on the contribution of each patch of the language branch to the alignment task for the UCF101 [33] dataset.
Figure 17: Statistics results on the contribution of each patch of the image branch to the alignment task for the UCF101 [33] dataset.
Figure 18: Statistics results on the contribution of each patch of the language branch to the alignment task for the Food101 [2] dataset.
Figure 18: Statistics results on the contribution of each patch of the language branch to the alignment task for the Food101 [2] dataset.
Figure 19: Statistics results on the contribution of each patch of the image branch to the alignment task for the Food101 [2] dataset.
Figure 20: Statistics results on the contribution of each patch of the language branch to the alignment task for the SUN397 [36] dataset.
Figure 20: Statistics results on the contribution of each patch of the language branch to the alignment task for the SUN397 [36] dataset.
Figure 21: Statistics results on the contribution of each patch of the image branch to the alignment task for the SUN397 [36] dataset.
Figure 22: Statistics results on the contribution of each patch of the language branch to the alignment task for the Caltech101 [11] dataset.
Figure 22: Statistics results on the contribution of each patch of the language branch to the alignment task for the Caltech101 [11] dataset.
Figure 23: Statistics results on the contribution of each patch of the image branch to the alignment task for the Caltech101 [11] dataset.

### 11 Mar 2024 v2

@@ Abstract
Prompt learning has emerged as an efficient alternative for fine-tuning foundational models, such as CLIP, for various downstream tasks. However, there is no work that provides a comprehensive explanation for the working mechanism of the multi-modal prompts. In this paper, we conduct a direct analysis of the multi-modal prompts by asking the following questions: $(i)$ How do the learned multi-modal prompts improve the recognition performance? $(ii)$ What do the multi-modal prompts learn? To answer these questions, we begin by isolating the component of the formula where the prompt influences the calculation of self-attention at each layer in two distinct ways, i.e., $(1)$ introducing prompt embeddings makes the $[cls]$ token focus on foreground objects. $(2)$ the prompts learn a bias term during the update of token embeddings, allowing the model to adapt to the target domain. Subsequently, we conduct extensive visualization and statistical experiments on the eleven diverse downstream recognition datasets. From the experiments, we reveal that the learned prompts improve the performance mainly through the second way, which acts as the dataset bias to improve the recognition performance of the pre-trained model on the corresponding dataset. Meanwhile, we propose the bias tuning way to validate our finding. With a deeper understanding of the multi-modal prompt, we hope our work can inspire new and solid research in this direction.
@@ 1 Introduction
In the alignment contribution statistic, we count the relevance between inputs and $[cls]$ token during alignment. In the statistics results of the text branch, we observe that the contribution of category tokens does not increase but decreases. The contribution of $t_{SOS}$ also decrease. On the contrary, the contribution of prompt tokens significantly increased. For the image branch, we reveal that the learned multi-modal prompts do not significantly adjust the relative relevance to $[cls]$ between inputs. Surprisingly, the contribution of multi-modal prompts to the alignment process is significant.
@@
To validate our findings, we propose a novel bias tuning that directly incorporates the learnable bias to each transformer block to validate the importance of the dataset bias. Comparative experiments demonstrate that bias tuning outperforms prompt tuning with the same number of parameters.
Main Findings and Contributions: (1) We probe to understand the working mechanism of the multi-modal prompts for the pre-trained vision-language model, conducting a series of extensive experiments from various perspectives, including alignment relevance, attention, attention distance, attention entropy, and visualization. Our findings show that the prompts mainly learn as the dataset bias. (2) We propose a novel bias tuning that directly incorporates the learnable bias to each transformer block to validate our finding.
2 Related Works
2.1 Vision Language models
2.2 Prompt learning
Prompt learning is an alternative fine-tuning method for transferring a model towards downstream tasks without re-learning the trained model parameters. This approach adapts a pre-trained model by adding a small number of new learnable embeddings at the input known as prompt tokens. Due to its efficiency in terms of parameters and convergence rate, prompt learning is found to be of great interest for adapting foundational models like CLIP for vision [17, 42, 34, 35] and vision-language tasks [44, 43, 46, 9]. CoOp [44] fine-tunes CLIP by optimizing a continuous set of prompt vectors in its language branch for few-shot image recognition. Bahng et al. [1] perform visual prompt tuning on CLIP by learning prompts on the vision branch. [6] and [22] propose to learn multiple sets of prompts for learning different contextual representations. CoCoOp [43] highlights the overfitting problem of CoOp and proposes to condition prompts based on visual features for improved performance on generalization tasks. MaPLe [18] proposes a multi-modal prompt learning approach by learning hierarchical prompts jointly at the vision and language branches of CLIP for better transfer. Our approach builds on the naturally independent prompt tuning way [31] where prompts are learned at both the vision and language encoder of CLIP.
@@ 3.1 Vision-Language Pre-trained Model (VLM)
Figure 2: Statistics results on the contribution of each token of the language branch to the alignment task for the OxfordPet dataset. We conduct the statistical analysis on samples where the zero-shot CLIP misidentify, while the independent prompt CLIP identify correctly. The corresponding two realizations are averaged over the statistics for the entire dataset, where the blue and red regions represent the statistics for zero-shot and prompt tuning, respectively. On the left side of the red vertical line, we show the contribution of $t_{SOS}$ and four learned text prompts $\bm{P}_{t}$. On the right side, we show the $\bm{c}_{k}$ tokens and padding tokens. The horizontal coordinate is the index of the token and the vertical coordinate is the relative value of the contribution. For the zero-shot clip, we use the temple: ‘‘a photo of [Category].’’. For the tokens after the $\bm{t}_{EOS}$, we set their contribution as 0 for intuitive comparison. Meanwhile, we zoom in on comparing the index at which the prompt is placed in the middle of each figure. The statistics results of other datasets are detailedly shown in the Appendix.
Figure 3: Statistics results on the contribution of each patch of the image branch to the alignment task for the OxfordPet dataset. We conduct the statistical analysis on samples where the zero-shot CLIP misidentify, while the independent prompt CLIP identify correctly. The corresponding two realizations are averaged over the statistics for the entire dataset, where the blue and red regions represent the statistics for zero-shot and prompt tuning, respectively. On the left side of the red vertical line, we show the contribution of the input patch tokens $\bm{\tilde{X}_{p}}$. On the right side, we show the learned vision prompts $\bm{P}_{v}$. The horizontal coordinate is the index of the token and the vertical coordinate is the relative value of the contribution. For the zero-shot clip, the values of the 196th-199th index are empty due to the absence of the vision prompts. Meanwhile, we zoom in on comparing the index at which the prompt is placed in the middle of each figure. The statistics results of other datasets are detailedly shown in the Appendix.
@@ 4.2.2 Aligning Contribution Statistics
To visually illustrate the dependency of alignment tasks on image patches and text tokens after incorporating prompts, we conduct a statistical analysis on eleven fine-grained recognition datasets. Before visualization, we preprocess the contribution according to normalization for intuitive comparison.
The distribution of relevance for the OxfordPet dataset is presented in Fig.2 and Fig.3.
  • In the statistics results of the text branch, we observe that the contribution of category tokens does not increase but decreases. The contribution of $t_{SOS}$ decreases, too. On the contrary, the contribution of the independent prompt tokens significantly increase.
@@ 4.3.1 Learned Textual prompts on Language Branch
Table 1: Validating the importance of bias. We compare bias with prompt tuning with existing methods in the same training settings.
Method Shot Setting
4 8 16
VPT 67.6 70.3 74.3
CoOP 69.4 72.9 76.6
IVLP 71.7 74.4 77.8
MaPLe 71.3 74.2 76.9
[…]
Ours 72.4 75.4 79.5
@@ 4.4 Validation for the importance of the bias
To validate the conclusion, we directly add the learnable bias to the pre-trained model and compare it with the prompt tuning. In the bias tuning setting, we add an independent learnable embedding and repeat it to add to all tokens after each transformer layer of the image and language branch. After learning in the same few-shot setting, we compare the bias tuning, prompt tuning, zero-shot CLIP, and some other existing methods in Table.1.
@@ 5 Conclusion
This paper probes to understand the working mechanism of the multi-modal prompts for the pre-trained vision-language model, conducting a series of extensive experiments from various perspectives, including alignment relevance, attention, attention distance, attention entropy, and visualization. Our findings show that the prompts mainly learn as the dataset bias. We propose a novel bias tuning that directly incorporates the learnable bias to each transformer block to validate our finding.